OpenStudy (anonymous):

Solve cos x tan x= 1/2 for x over the interval [0,2Pi] A.) Pi/6 B.) Pi/3 and 5Pi/3 C.) Pi/6 and 11Pi/6 D.) Pi/6 and 5Pi/6

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \tan x=\frac{ \sin x }{\cos x} }\) therefore, \(\color{#000000 }{ \displaystyle \cos x \tan x=\frac{\cos x \sin x }{\cos x} =\frac{\bcancel{\cos x} \sin x }{\bcancel{\cos x}} }\)

OpenStudy (solomonzelman):

So, cosines are going to be cancelled \(....\) However, you'll have to exclude any solutions for which \(\color{black}{\cos x=0}\).

OpenStudy (solomonzelman):

(not that any solutions will have \(\color{black}{\cos x=0}\))

OpenStudy (anonymous):

I'm still kinda lost. Is it B- \[\frac{ \Pi }{ 3 } and \frac{ 5\Pi }{ 3 }\]

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@roadjester - Is my answer correct? Can you help me, please? Im horrible with Algebra II/ Trig

OpenStudy (solomonzelman):

Please, show your work that allowed you to arrive at yout answer.

jhonyy9 (jhonyy9):

@TheMadden5 do you know the answer ?

jhonyy9 (jhonyy9):

so than you know that sin30° = 1/2 - what will be the right answer ?

OpenStudy (anonymous):

Ive figured it out. Thanks for helping everyone!